# Generative Image as Action Models

### v1

@@ 1 Introduction
In this work, we use image-generation models in their native formulation: drawing images. We present Genima, a multi-task behavior-cloning agent that directly fine-tunes Stable Diffusion [1] to “draw joint-actions”. To supervise the fine-tuning, we format expert demonstrations into an image-to-image dataset. The input is an RGB image with a language goal, and the output is the same image with joint-position targets from a future timestep in the demonstration. The targets are rendered as colored spheres for each joint, as shown on the previous page. These visual targets are fed into a controller that maps them to a sequence of joint-positions. This formulation frames action-generation as an image-generation problem such that action-patterns become visual-patterns.
@@ 2 Genima
Genima is a behavior-cloning agent that maps RGB observations and a language goal g into joint-position actions . The key objective is to lift actions into image-space such that internet-pretrained diffusion models can learn action-patterns as visual-patterns. We achieve this with a two-stage process: fine-tune Stable Diffusion [1] to draw images with target joint-positions that are timesteps ahead, and train a controller to translate these targets into to a sequence of executable joint-positions . This simple two-stage process offloads semantic and task-level reasoning to a generalist image-generation model, while the controller reaches nearby joint-positions indicated by the visual targets. The sections below describe the two stages in detail, and Figure 1 provides an overview.
@@ 2.1 Diffusion Agent
The diffusion agent controls what to do next. The agent takes RGB observations and language goal g as input, and outputs an image with target joint-positions . This problem formulation is a classic image-to-image generation setup, so any diffusion fine-tuning pipeline [7, 28, 29] can be used. We specifically use ControlNet [8] to preserve spatial layouts and for data-efficient fine-tuning.
Fine-Tuning Data. To supervise the fine-tuning, we take expert demonstrations, and randomly sample observations and target joint-positions from timesteps. For that timestep, we obtain 6-DoF poses of each robot joint, which is available through robot-APIs (from forward-kinematics). We place spheres at those poses with pyrender111https://pyrender.readthedocs.io/en/latest/examples/quickstart.html, and render them on four camera observations with known intrinsics and extrinsics. On a 7-DoF Franka Panda, we only render four joints: base, elbow, wrist, and gripper, to avoid cluttering the image. Each joint is represented with an identifying color, with separate colors for gripper open and close. The spheres include horizontal stripes parallel to the joint’s rotation axis, acting as graduations indicating the degree of rotation. Only horizontal stripes are needed as each joint has only one rotation axis.
Fine-Tuning with ControlNet. Given an image-to-image dataset, we finetune Stable Diffusion [27] with ControlNet [8] to draw targets on observations. ControlNet is a two-stream architecture: one stream with a frozen Stable Diffusion UNet that gets noisy input and language descriptions, and a trainable second stream that gets a conditioning image to modulate the output. This architecture retains the text-to-image capabilities of Stable Diffusion, while fine-tuning outputs to spatial layouts in the conditioning image. Genima uses RGB observations as the conditioning image to draw precise targets. We use SD-Turbo [27] – a distilled model that can generate high-quality images within 1 to 4 diffusion steps – as the base model for fine-tuning. We use the HuggingFace implementation [30]222https://huggingface.co/docs/diffusers/en/using-diffusers/controlnet of ControlNet without modifications. See Appendix D for more details on fine-tuning.
@@ Acknowledgments
Big thanks to the members of the Dyson Robot Learning Lab for discussions and infrastructure help: Nic Backshall, Nikita Chernyadev, Iain Haughton, Yunfan Lu, Xiao Ma, Sumit Patidar, Younggyo Seo, Sridhar Sola, Eugene Teoh, Jafar Uruç, and Vitalis Vosylius. Special thanks to Tony Z. Zhao and Cheng Chi for open-sourcing ACT and Diffusion Policies, and the HuggingFace Team for diffusers.

### v2

@@ 1 Introduction
In this work, we use image-generation models in their native formulation: drawing images. We present Genima, a multi-task behavior-cloning agent that directly fine-tunes Stable Diffusion [1] to “draw joint-actions”. To supervise the fine-tuning, we format expert demonstrations into an image-to-image dataset. The input is an RGB image with a language goal, and the output is the same image with joint-position targets from a future timestep in the demonstration. The targets are rendered as colored spheres for each joint (as shown in the previous page). These visual targets are fed into a controller that maps them to a sequence of joint-positions. This formulation frames action-generation as an image-generation problem such that action-patterns become visual-patterns.
@@ 2 Genima
Genima is a behavior-cloning agent that maps RGB observations and a language goal g to joint-position actions . The key objective is to lift actions into image-space such that internet-pretrained diffusion models can learn action-patterns as visual-patterns. This is accomplished through a two-stage process: fine-tuning Stable Diffusion [1] to draw target joint-positions on input images, timesteps ahead, and training a controller to translate these targets into to a sequence of executable joint-positions . This simple two-stage process offloads semantic and task-level reasoning to a generalist image-generation model, while the controller reaches nearby joint-positions indicated by the visual targets. The sections below describe the two stages in detail, and Figure 1 provides an overview.
@@ 2.1 Diffusion Agent
The diffusion agent controls what to do next. The agent takes language goal g and multi-view RGB images as input, and outputs target joint-positions on the same images. This problem formulation is a classic image-to-image setting, so any fine-tuning pipeline [7, 28, 29] can be used. We specifically use ControlNet [8] to preserve spatial-layouts and for data-efficient fine-tuning.
Fine-Tuning Data. To supervise the fine-tuning, we randomly sample observations and target joint-positions from timesteps in expert demonstrations. For that timestep, we obtain 6-DoF poses of each robot joint, which is available through robot-APIs (from forward-kinematics). We place spheres at those poses with pyrender111https://pyrender.readthedocs.io/en/latest/examples/quickstart.html, and render them on four camera observations with known intrinsics and extrinsics. On a 7-DoF Franka Panda, we only render four joints: base, elbow, wrist, and gripper, to avoid cluttering the image. Each joint is represented with an identifying color, with separate colors for gripper open and close. The spheres include horizontal stripes parallel to the joint’s rotation axis, acting as graduations indicating the degree of rotation. Only horizontal stripes are necessary as each joint has only one rotation axis. The stripes are also asymmetric across the poles to help identify the sphere orientation.
Fine-Tuning with ControlNet. Given an image-to-image dataset, we finetune Stable Diffusion [27] with ControlNet [8] to draw targets on observation images. ControlNet is a two-stream architecture: one stream with a frozen Stable Diffusion UNet that gets noisy input and language descriptions, and a trainable second stream that gets a conditioning image to modulate the output. This architecture retains the text-to-image capabilities of Stable Diffusion, while fine-tuning outputs to spatial layouts in the conditioning image. Genima uses RGB observations as the conditioning image to draw precise targets. We use SD-Turbo [27] – a distilled model that can generate high-quality images within 1 to 4 diffusion steps – as the base model for fine-tuning. We use the HuggingFace implementation [30]222https://huggingface.co/docs/diffusers/en/using-diffusers/controlnet of ControlNet without modifications. See Appendix D for more details on fine-tuning.
@@ Acknowledgments
Big thanks to the members of the Dyson Robot Learning Lab for discussions and infrastructure help: Abdi Abdinur, Nic Backshall, Nikita Chernyadev, Iain Haughton, Yunfan Lu, Xiao Ma, Sumit Patidar, Younggyo Seo, Sridhar Sola, Eugene Teoh, Jafar Uruç, and Vitalis Vosylius. Special thanks to Tony Z. Zhao and Cheng Chi for open-sourcing ACT and Diffusion Policies, and the HuggingFace Team for diffusers.
@@ Appendix I SuSIE Goal Predictions
Hallucinations in predictions. As with any image-generation framework, Genima’s diffusion agent is prone to hallucinating visual artifacts. These hallucinations are limited to joint-action spheres since the background observations remain unchanged. A potential solution could be to use a classifier to detect out-of-distribution actions and then execute recovery behaviors.